OpenStudy (anonymous):

A parallelogram with sides of 6 and 10 has an area of 30 sq root of 2. Find the measure of the small angle of the parallelogram. 30 45 60

OpenStudy (anonymous):

base*height = area\[10 h = 30 \sqrt{2}\]\[h\text{=}3 \sqrt{2}\]\[\theta =\text{ArcSin}\left[\frac{3 \sqrt{2}}{6}\right]180/\pi \text{ = }45{}^{\circ} \]